OpenStudy (anonymous):

If a sample of barium contains 2.3 x 1023 atoms of barium, how many grams of barium are in the sample? 0.38 g 2.62 g 52.4 g 359.5 g

OpenStudy (jfraser):

What is the relationship between the # of particles and their mass? Should be a big ##

OpenStudy (anonymous):

n(number of moles)=2.3*10^23/6.02*10^23[known as Avogadro constant]=0.381 mol so the mass is m=n*N=0.381mol*137g/mol

OpenStudy (anonymous):

it's about 52.4